OpenStudy (anonymous):

is -6(2x-10)+12< or = to 180 sometimes, always, or never true?

OpenStudy (cwrw238):

-6(2x-10)+12 = -12x + 60 + 12 = -12x + 72 so we have -12x + 72 <= 180

OpenStudy (cwrw238):

or -12x <= 108

OpenStudy (cwrw238):

so consider if x is a high negative value e.g. - 20 does that fit the inequality?

OpenStudy (anonymous):

can you help me? how do you solve this: 6[5y-(3y-1)] . or = to 4(3y-7) thanks for your help on the other question.

OpenStudy (cwrw238):

wheres the inequality sign?

OpenStudy (anonymous):

oops.. it's this: > or =to

OpenStudy (precal):

 is this it?